A bus traveled on a level road for 6 hours at an average speed 20mph faster that it traveled on a winding road. The time spent on the winding road was 2 hours. Find the average speed on the level road is the entire trip was 504 miles.
Answer by mananth(16946)

speed on level rad = x
then speed on winding rod be (x-20)
d=rt
6x+2(x-20)=504
6x +2x -40 = 504
8x-40 = 504
8x=504
x= 63
speed on level road is 63 mph and on winding road = 43 mph